A highway connects city "A" and city "B" and extends 2659 miles. Paul Simon begins in city "A" and drives at an average rate of 43 MPH on the highway toward city B. At the same time John Smith begins in city "B" and drives on the highway at an average rate of 55 MPH toward city "A". If Paul and John maintain these average speeds, how long after they start will they meet?
___hr (Round to the nearest hundredth as needed)

Answer by stanbon(75887)

A highway connects city "A" and city "B" and extends 2659 miles. Paul Simon begins in city "A" and drives at an average rate of 43 MPH on the highway toward city B. At the same time John Smith begins in city "B" and drives on the highway at an average rate of 55 MPH toward city "A". If Paul and John maintain these average speeds, how long after they start will they meet?
-----------
Their approaching rate is 43+55 = 98 mph
-------------------
Equation:
time = distance/rate = 2659/98 = 27.13 hrs = 27 hr 8 minutes
========================================
Cheers,
Stan H.
========================================